OpenStudy (anonymous):

Solve the equation Quantity x minus 5 all over 4 = 3 for x. Please show your work

OpenStudy (anonymous):

x - 5 = 12 x = 17

OpenStudy (lilg132):

\[{x - 5 \over 4} = 3\]

OpenStudy (anonymous):

lilg sorry but this is incorrect

OpenStudy (lilg132):

how did you work it out?

OpenStudy (anonymous):

multiplies four on both sides x - 5 = 12 x = 17

OpenStudy (anonymous):

(x-5)/4 = 3 eliminate denominator 4 by multiplying both sides of the equation with 4 4[(x-5)/4)] = 3(4) x-5 = 12 solve for x by adding 5 on both sides of the equation 5+x-5 = 12+5 x=17